prove that the element (u,v) in direct product r×s of two rings r and s is unit if and only if u is a unit in r and v is unit in s


1
Expert's answer
2021-02-24T14:22:31-0500

Let's first show, that "(u,v)" is an identity element. Indeed, by definition of ring structure on a product "R\times S" we have:

"(u,v)\cdot(r,s)=(u\cdot r,v\cdot s)=(r,s)"

"(r,s)\cdot(u,v)=(r\cdot u,s\cdot v)=(r,s)"

Now as identity is unique, the identity element on "R\times S" should forcefully be "(u,v)".